OpenStudy (anonymous):

Solve this system of equations. Write the answer as an ordered pair. -5x - 2y = -2 x - 7y = -7 ....... i think i have the right answer but not sure, can somebody reassure me

OpenStudy (uri):

@kyle_bradley Can you please show your work?

OpenStudy (anonymous):

the ans u came up with plug them in and if u get the statement true then u came up with the correct ans

OpenStudy (anonymous):

yes, -5(7y=7)-2y=-2....and i just realized that it didnt ask for the equation

OpenStudy (anonymous):

oops, its -5(7y-7)-2y=-2

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@MaimiGirl